### v1

@@ 5 Symmetric investment games
We now apply our theoretical framework on some games for which implementing reaction functions through blockchain-based smart contracts seems very promising. These are symmetric $n$-player games with actions $A_{i}=\{0,\dots,H\}$ interpreted as investment amounts. Reaction functions now go from being a theoretical concept to a most practical tool: conditional investment strategies provide an excellent use case for smart contracts (e.g. Weber and Staples, 2022). Payoffs for our symmetric investment games include a non-decreasing common value $v\colon A\to\mathbb{R}$ derived symmetrically from costly individual actions. [^6] That is, for each outcome $a\in A$ and permutation $\pi\colon N\to N$, we have $v(a)=v(a_{\pi(1)},\dots,a_{\pi(n)})$. Specifically, for each player $i$, and outcome $a\in A$,
@@
As players are symmetric, coordinated outcomes ($a_{i}=a_{j}$ for all players $i$ and $j$) will be key. For convenience, we assume that these coordinated outcomes are Pareto ranked and that the higher players coordinate the better: $u_{i}(\alpha,\dots,\alpha)$ is increasing in $\alpha\in\{0,\dots,H\}$. [^7] This can be relaxed and we instead obtain a variation on Theorem 5.2. Additional structure will be imposed on $v$ in Subsections 5.1 and 5.2, resulting in weakest link and public good games, respectively.
In what follows, we focus on monotone reaction functions (“invest more the more others invest”). Monotone reaction functions make fixed-point existence and computation straightforward. Let $\mathcal{M}_{i}\subseteq\mathcal{R}_{i}$ denote player $i$’s monotone reaction functions.
Moreover, as action spaces $A_{i}$ are identical and contribution to the total value $v$ is symmetric across players, it is natural to further constrain to symmetric reaction functions. In essence, the identities of the other players should not matter. A reaction function $R_{i}\in\mathcal{R}_{i}$ is symmetric if, for each $a_{-i}\in A_{-i}$ and permutation $\pi\colon N\to N$ with $\pi(i)=i$, we have $R_{i}(a_{-i})=R_{i}(\pi(a)_{-i})$ with $\pi(a)\equiv(a_{\pi(1)},\dots,a_{\pi(n)})\in A$. Let $\mathcal{S}_{i}\subseteq\mathcal{M}_{i}$ denote player $i$’s symmetric and monotone reaction functions. [^8] For instance, in oligopoly models, a firm’s profit typically is unchanged if its competitors’ actions (prices, quantities) are permuted. In this case, a symmetric reaction function specifies the same reaction to two such permuted action profiles.
With symmetric reaction functions, it becomes meaningful to compare functions across players. We say that $i$ and $j$ play the same reaction function, denoted $R_{i}\cong R_{j}$, whenever $R_{i}(x)=R_{j}(x)$ for each $x\in\{1,\dots,H\}^{n-1}$. Let $\mathcal{N}\subset\mathcal{S}\subset\mathcal{R}$ be profiles (“social norms”) in which all play the same, that is, $R_{i}\cong R_{j}$ for all players $i$ and $j$. A reaction function $R^{*}_{i}$ is norm-proof if, for every norm $R\in\mathcal{N}$, playing $R^{*}_{i}$ against the norm $R$ yields at least as high payoff as conforming to it. This is not an equilibrium property, in the sense that we do not require $R^{*}_{i}$ to be optimal against a specific $R_{-i}$, but rather an alternative interpretation of safe play. Even if you do not know which norm $R$ the others might settle on, playing $R^{*}_{i}$ is a safe and reasonable choice at least as good as conforming to the norm.
@@
Reaction function $R^{*}_{i}\in\mathcal{M}_{i}$ is norm-proof if, for each profile $R\in\mathcal{N}$,
@@
Technically, the property bears resemblance to the “evolutionary stable strategy” (ESS) refinement of Nash equilibrium prominent in evolutionary game theory (Maynard Smith and Price, 1973). [^9] Further evidence to this point is given in the Appendix, where we simulate “evolution” of a pool of reaction functions in a public-good setting. The simulations indicate that evolutionary selection indeed would push towards a pool of mainly norm-proof reactions.
Norm-proof reaction functions have a certain structure in symmetric investment games. We say that a reaction function $R_{i}$ exhibits conditional collaboration if it satisfies the following two criteria: if the others coordinate, then $i$ does as well; if not, then $i$ matches at least the minimum investment but undercuts the maximum investment. That is, $R_{i}\in\mathcal{R}_{i}$ exhibits conditional collaboration if, for each $\alpha\in\{1,\dots,H\}$, $R_{i}(\alpha,\dots,\alpha)=\alpha$ and, for all remaining $a_{-i}\in A_{-i}$,
$\min_{j\neq i}a_{j}\leq R_{i}(a_{-i})<\max_{j\neq i}a_{j}.$
The principle of conditional collaboration is familiar from behavioral economics, where experimental evidence has shown that participants are often willing to contribute more the more others contribute (e.g. Fischbacher et al., 2001). This has typically been rationalized by preferences reflecting altruism, “warm glow”, or reciprocity (e.g. Sugden, 1984; Croson, 2007). We offer an alternative perspective where conditional collaboration arises even without such “moral preferences”. For reaction functions, conditional collaboration can instead be rationalized through norm-proofness.
Theorem 5.1.
@@
There is a unique reaction function exhibiting conditional collaboration if there are only two players (“match the other’s action”) or only two ordered actions (“match the minimum action of the others”), both which are safe. However, with more players and actions, the class is large (and may contain unsafe reactions). In common to all of them is that, for symmetric investment games, the desirable high-investment outcome is always supported in reaction-function equilibrium.
Theorem 5.2.
For symmetric investment games, all profiles consisting only of reaction functions that exhibit conditional collaboration support the high-investment outcome $(H,\dots,H)$ in reaction-function equilibrium.
@@
It is immediate that payoff-consistent reaction functions exist (e.g., constant reactions). The best-reply reaction (with a fixed tie-breaking rule if needed) is also payoff consistent. In many games, the property even has no bite whatsoever as there might not be any payoff-equivalent action profiles. However, in the games that we explore next in Subsections 5.1 and 5.2, payoff consistency combines with norm-proofness to give a unique recommendation.
@@ 5.1 Weakest-link games
In weakest-link (or “minimum-effort”) games, the minimum investment determines the outcome; the higher this is, the better for all (e.g. Hirshleifer, 1983). These games are abundant in practice and arise, for instance, when divisions within a firm contribute independently to a joint project for which the lowest-quality input disproportionally affects the output (Knez and Camerer, 1994; Kunreuther, 2009; Dietrichson et al., 2022). Specifically, payoffs are given by
@@
with parameter $\lambda>1$. Hence, players wish to match the minimum investment of the others and ideally do so at $H$. Safe play amounts to taking action $a_{i}=0$ with $\underaccent{\bar}{v}_{i}=u_{i}(0,\dots,0)=0$. Here, action profiles $a_{-i}$ and $a^{\prime}_{-i}$ are payoff equivalent whenever $\min_{j\neq i}a_{j}=\min_{j\neq i}a^{\prime}_{j}$.
The best-reply reaction $\textit{BR}_{i}$ is particularly simple for weakest-link games and entails matching the minimum investment, $\textit{BR}_{i}(a_{-i})=\min_{j\neq i}a_{j}$. Not only is it safe, it is also the only norm-proof and payoff-consistent reaction function. [^10] Combining proofs, one may also show that $\textit{BR}_{i}$ is the only norm-proof reaction that is safe for all parameter values $\lambda>1$.
@@
In weakest-link games, there are compelling reasons to use monotone reaction functions. If we also can expect others to use such functions, then we obtain a different argument in favor of $\textit{BR}_{i}$. For this purpose, we restrict to what we call high-risk weakest-link games with parameter $\lambda<H/(H-1)$. That is, coordination gains are small relative to effort costs to the point that coordinating at 00 is superior to playing $H$ against $H-1$. This does not mean that there is no gain from coordinating at a higher investment level—the payoff is still higher at $(H,\dots,H)$ than at $(0,\dots,0)$—rather, it means that higher investments comes with greater risk. [^11] For comparison, when Van Huyck et al. (1990) found that experiment participants failed to coordinate, the setting corresponded to $H=6$ and $\lambda=2$. This is a much easier environment to coordinate in than high-risk games, where we would have $\lambda<H/(H-1)=1.2$.
@@
The high-risk qualification is restrictive, but this is to be expected. We already know from Theorem 5.2 that many reaction functions “solve” weakest-link games in general. The purpose of Theorem 5.4 is partly to identify a strong-enough condition under which we can make a unique, precise recommendation.
@@
Consider the two-player weakest-link game. Suppose, for contradiction, that reaction function $\hat{R}_{i}\in\mathcal{R}_{i}$ is weakly dominant. Suppose first the other player chooses $R_{j}$, which is to always play $H$. Then we must have $\hat{R}_{i}(H)=H$ to ensure that the optimal outcome $(H,H)$ is a fixed point. Consider next the case in which the other player chooses reaction $R^{\prime}_{j}$ such that $R^{\prime}_{j}(H-1)=H$ and otherwise $R^{\prime}_{j}(a_{i})=0$, which is neither monotone nor safe (in a high-risk setting). The best attainable fixed point now is $(a_{i},a_{j})=(H-1,H)$, but this would require $\hat{R}_{i}(H)=H-1$, which is a contradiction. Hence, no reaction function is optimal against both $R_{j}$ and $R^{\prime}_{j}$. $\circ$
The fundamental message to take from the above is that reaction functions offer a new way to overcome trust issues in weakest-link games. Reaction-function play does not rely on building trust between the players (Chen and Chen, 2011), on putting hope to communication routines (Blume and Ortmann, 2007; Dietrichson et al., 2022), or on establishing a central authority to monitor and punish inefficient action (Holmstrom, 1982). Neither of this is needed—the solution would in principle allow any large group of strangers to coordinate on their very first attempt.
Once we dig deeper and identify which particular reaction function to play, it is no surprise that the best-reply reaction function stands out (i.e., to match the weakest link). In a sense, Theorems 5.3 and 5.4 together with the surrounding discussion make for a robustness exercise that verifies that, indeed, the intuitive best-reply reaction is the best option.
@@ 5.2 Public-good games
We turn to a different class of games for which, in contrast, best replies are ineffective. We have already seen that reaction functions introduce new ways for players to coordinate around Pareto-improving outcomes in strictly dominated actions (e.g. in the Prisoner’s dilemma). Now, we will further expand on this in the particular context of public-good games, in which each player’s investment benefits also all others, but the investment is a net loss to the player herself (e.g. Bergstrom et al., 1986). In public-good games, payoffs are given by
@@
with parameter $11/n<\lambda<$ 1.121212Outside this parameter range, high investment (for $\lambda\geq 1$) or zero investment (for $\lambda\leq 1/n$) is both safe, dominant, and efficient. High investment $(H,\dots,H)$ maximizes total payoffs $\sum_{j}u_{j}(a)$ but, in the strategic game, free-riding through $a_{i}=0$ is dominant. Hence, the best reply is simply to free-ride and never invest. Again, safe play yields $\underaccent{\bar}{v}_{i}=u_{i}(0,\dots,0)=0$. Now, action profiles $a_{-i}$ and $a^{\prime}_{-i}$ are payoff equivalent if they add up to the same: $\sum_{j\neq i}a_{j}=\sum_{j\neq i}a^{\prime}_{j}$.
[…]
@@
For public-good games, reaction function $R_{i}\in\mathcal{R}_{i}$ is norm-proof and payoff consistent if and only if, for each $a_{-i}\in A_{-i}$,
$R_{i}(a_{-i})=R^{*}_{i}(a_{-i})\equiv\left\lfloor\frac{1}{n-1}\textstyle\sum_{j\neq i}a_{j}\right\rfloor.$
One might overlook the subtleties associated to rounding down and propose a different reaction function, such as “match the average, rounded up”. By our definition, this does not exhibit conditional collaboration. A profile of such reaction functions is not an equilibrium—for instance, players may be able to deviate even by free-riding (specifically, if $n>H+1$). In this way, rounding down is important to sustain an equilibrium.
In the experimental literature, some attention has been given social-welfare preferences (Charness and Rabin, 2002), where a player’s preference depends also on the payoffs of others. It is possible to interpret the reaction $R^{*}_{i}$ in a similar way. Specifically, we define a reaction function $R_{i}\in\mathcal{R}_{i}$ to be welfare-maximizing conditional on safe play if, for each $a_{-i}\in A_{-i}$,
@@
As a parallel to Theorem 5.4, we define high-risk public-good games with parameters $\lambda<H/(nH-1)$. The intuition is the same: there are still gains from coordinating on higher investment, but the returns are small compared to the risks involved in overinvesting relative to the others. For these games, reaction $R^{*}_{i}$ is welfare maximizing subject to safe play. [^14] For weakest-link games, this property instead singles out $\textit{BR}_{i}$. There is no reason for $i$ to exceed the minimum investment as this creates no added value to the others but only adds to $i$’s investment costs; it is also not beneficial for $i$ to undercut the minimum, because this harms all players more than it saves on $i$’s costs.
@@
For high-risk public-good games, $R^{*}_{i}$ is welfare-maximizing conditional on safe play.
In this way, reaction functions provide new ways to overcome tendencies to free-ride in public-goods games, which does not rely on, for instance, players forming sanctioning institutions (Kosfeld et al., 2009), being altruistic (Anderson et al., 1998), or the introduction of obligations (Galbiati and Vertova, 2008).
Our main candidate for public-good games $R^{*}_{i}$ is to match the average, rounded down. We do not claim that this always is optimal (e.g., it is better to free-ride if the others unconditionally contribute), but it appears always quite reasonable: it is simple, safe, and ensures that the player is no worse off than the average. If faced with a pool of free-riders, it specifies to also not invest; against both conditional and unconditional collaborators, the high-investment outcome is realized. In this way, it does reasonably well against most reaction functions. This is further corroborated by the simulation study in the Appendix. There, in a specific public-good setting, we find compelling evidence that a process of evolutionary selection would converge towards functions alike $R^{*}_{i}$.

### v2

@@ 5 Symmetric investment games
We now apply our theoretical framework on some games for which implementing reaction functions through blockchain-based smart contracts seems very promising. These are symmetric $n$-player games with finite action sets $A_{i}=\{0,1,\dots,H\}$ interpreted as investment amounts. Reaction functions now go from being a theoretical concept to a most practical tool. Conditional investment strategies provide an excellent use case for smart contracts (e.g. Weber and Staples, 2022). Payoffs for symmetric investment games include a non-decreasing common value $v\colon A\to\mathbb{R}$ derived symmetrically from costly individual actions. [^8] That is, for each outcome $a\in A$ and permutation $\pi\colon N\to N$, we have $v(a)=v(a_{\pi(1)},\dots,a_{\pi(n)})$. Specifically, for each player $i$, and outcome $a\in A$,
@@
As players are symmetric, coordinated outcomes ($a_{i}=a_{j}$ for all players $i$ and $j$) will be key. For convenience, we assume that these coordinated outcomes are Pareto ranked and that the higher players coordinate the better: $u_{i}(\alpha,\dots,\alpha)$ is increasing in $\alpha\in\{0,\dots,H\}$. Additional structure will be imposed on $v$ resulting in weakest-link (Subsection 5.1) and public-good games (Subsection 5.2), respectively.
In what follows, we focus on monotone reaction functions (i.e., players commit to invest more the more others invest). Monotone reaction functions guarantee fixed-point existence and makes computation of such fixed points easy. Let $\mathcal{M}_{i}\subseteq\mathcal{R}_{i}$ denote the set of player $i$’s monotone reaction functions.
As action spaces $A_{i}$ are identical and contribution to the total value $v$ is symmetric across players, it is natural to consider symmetric reaction functions. In essence, the identities of the other players should not matter. A reaction function $R_{i}\in\mathcal{R}_{i}$ is symmetric if, for each $a_{-i}\in A_{-i}$ and permutation $\pi\colon N\to N$ with $\pi(i)=i$, we have $R_{i}(a_{-i})=R_{i}(a_{\pi(-i)})$ with $a_{\pi}\equiv(a_{\pi(1)},\dots,a_{\pi(n)})\in A$. Let $\mathcal{S}_{i}\subseteq\mathcal{M}_{i}$ denote the set of player $i$’s symmetric and monotone reaction functions. [^9] For instance, in oligopoly models, a firm’s profit typically is unchanged if its competitors’ actions (prices, quantities) are permuted. In this case, a symmetric reaction function specifies the same reaction to two such permuted action profiles.
With symmetric reaction functions, it becomes meaningful to compare functions across players. We say that $i$ and $j$ play the same reaction function, denoted $R_{i}\cong R_{j}$, whenever $R_{i}(x)=R_{j}(x)$ for each $x\in\{0,\dots,H\}^{n-1}$. That is, players $i$ and $j$ react in the same way to the actions of others. Let $\mathcal{N}\subset\mathcal{S}\subset\mathcal{R}$ be the set of profiles in which all players react in the same way, that is, $R_{i}\cong R_{j}$ for all players $i$ and $j$. We think of such profiles as social norms or conventions, and say that a reaction function $R^{*}_{i}$ is norm-proof if, for every norm $R\in\mathcal{N}$, playing $R^{*}_{i}$ against the norm $R$ yields at least as high payoff as conforming to it. [^10] Technically, the property bears resemblance to the “evolutionary stable strategy” (ESS) refinement of Nash equilibrium prominent in evolutionary game theory (Maynard Smith and Price, 1973). Further evidence to this point is given in the Appendix, where we simulate “evolution” of a pool of reaction functions in a public-good setting. The simulations indicate that evolutionary selection indeed would push towards a pool of mainly norm-proof reactions.
@@
Reaction function $R^{*}_{i}\in\mathcal{M}_{i}$ is norm-proof if, for each norm $R\in\mathcal{N}$,
@@
This is not an equilibrium property, in the sense that we do not require $R^{*}_{i}$ to be optimal against a specific $R_{-i}$, but rather an alternative interpretation of safe play. Even if you do not know which norm $R$ the others might settle on, playing $R^{*}_{i}$ is a safe and reasonable choice at least as good as conforming to the norm.
Norm-proof reaction functions turn out to have a certain structure in symmetric investment games. We say that a reaction function $R_{i}$ exhibits conditional collaboration if it satisfies the following two criteria: if the others coordinate (play the same action), then $i$ does as well; if not, then $i$ matches at least the minimum investment and at most the maximum investment. That is, $R_{i}\in\mathcal{R}_{i}$ exhibits conditional collaboration if, for each $\alpha\in\{0,\dots,H\}$, $R_{i}(\alpha,\dots,\alpha)=\alpha$ and, for all remaining $a_{-i}\in A_{-i}$,
$\min_{j\neq i}a_{j}\leq R_{i}(a_{-i})\leq\max_{j\neq i}a_{j}.$
The principle of conditional collaboration is familiar from behavioral economics, where experimental evidence has shown that participants are often willing to contribute more the more others contribute (e.g. Fischbacher et al., 2001). This has typically been rationalized by preferences reflecting altruism or reciprocity (e.g. Sugden, 1984; Croson, 2007). We offer an alternative perspective where conditional collaboration arises even without such “moral preferences”. For reaction functions, conditional collaboration can be rationalized through norm-proofness.
Observation 3.
@@
There is a unique reaction function exhibiting conditional collaboration if there are only two players (“match the other’s action”), which is safe. However, with more players, the class is large and may contain unsafe reactions.
@@
It is immediate that payoff-consistent reaction functions exist (e.g., constant reactions). The best-reply reaction (with a fixed tie-breaking rule if needed) is also payoff consistent. For many games, the property will have no bite whatsoever as there might not be any payoff-equivalent action profiles. However, in the games that we explore next in Subsections 5.1 and 5.2, payoff consistency combines with norm-proofness to give a unique recommendation.
@@ 5.1 Weakest-link games
In weakest-link (or “minimum-effort”) games, the minimum investment determines the players’ common value $v(a)$; the higher this is, the better for all (e.g. Hirshleifer, 1983). These games are abundant in practice and arise, for instance, when divisions within a firm contribute independently to a joint project for which the lowest-quality input disproportionally affects the output (Knez and Camerer, 1994; Kunreuther, 2009; Dietrichson et al., 2022). Specifically, payoffs are given by
@@
with parameter $\lambda>1$. Hence, players wish to match the minimum investment of the others and ideally do so at $H$. Conventional safe play amounts to taking action $a_{i}=0$ with $\underaccent{\bar}{v}_{i}=u_{i}(0,\dots,0)=0$. Moreover, action profiles $a_{-i}$ and $a^{\prime}_{-i}$ are payoff equivalent whenever $\min_{j\neq i}a_{j}=\min_{j\neq i}a^{\prime}_{j}$.
The best-reply reaction $\textit{BR}_{i}$ is particularly simple for weakest-link games and entails matching the minimum investment, $\textit{BR}_{i}(a_{-i})=\min_{j\neq i}a_{j}$. Not only is it safe, it is also the only norm-proof and payoff-consistent reaction function. [^11] In fact, $\textit{BR}_{i}$ is the only norm-proof reaction that is safe for all parameter values $\lambda>1$.
@@
Playing $BR_{i}$ against a profile of reaction functions that exhibit conditional collaboration always supports the high-investment outcome $(H,\dots,H)$ in reaction-function equilibrium.
In weakest-link games, there are compelling reasons to expect that players use monotone reaction functions, in which case we obtain an additional argument in favor of $\textit{BR}_{i}$. For this, we restrict attention to what we call high-risk weakest-link games with parameter $\lambda$ such that $1<\lambda<H/(H-1)$. That is, coordination gains are small relative to effort costs to the point that coordinating at 0 is superior to playing $H$ against H1H-1. This does not mean that there is no gain from coordinating at a higher investment level—the payoff is still higher at $(H,\dots,H)$ than at $(0,\dots,0)$—rather, it means that higher investments come with greater risk. [^12] For comparison, when Van Huyck et al. (1990) found that experiment participants failed to coordinate, the setting corresponded to $H=6$ and $\lambda=2$. This is a much easier environment to coordinate in than high-risk games, where we would have $\lambda<H/(H-1)=1.2$.
@@
The high-risk qualification is restrictive, but this is to be expected. There are many reaction functions that “solve” weakest-link games in general. The purpose of Theorem 5.2 is partly to identify a strong-enough condition under which we can make a unique, precise recommendation of which reaction function to play.
@@
Consider a two-player weakest-link game. Suppose, for contradiction, that there exists a reaction function $\hat{R}_{i}\in\mathcal{R}_{i}$ that is weakly dominant in general (so not only against monotone play, for instance). Suppose first the other player chooses $R_{j}$, which is to always play $H$. Then we must have $\hat{R}_{i}(H)=H$ to ensure that the optimal outcome $(H,H)$ is a fixed point. Consider next the case in which the other player chooses reaction $R^{\prime}_{j}$ such that $R^{\prime}_{j}(H-1)=H$ and otherwise $R^{\prime}_{j}(a_{i})=0$, which is neither monotone nor safe (in a high-risk setting). The best attainable fixed point now is $(a_{i},a_{j})=(H-1,H)$, but this would require $\hat{R}_{i}(H)=H-1$, which is a contradiction. Hence, no reaction function is optimal against both $R_{j}$ and $R^{\prime}_{j}$. $\circ$
The fundamental message to take from this section is that reaction functions offer a new way to overcome trust issues in weakest-link games. Reaction-function play does not rely on building trust between the players (Chen and Chen, 2011), on putting hope to communication routines (Blume and Ortmann, 2007; Dietrichson et al., 2022), or on establishing a central authority to monitor and punish inefficient action (Holmstrom, 1982). None of this is needed—using reaction functions as commitment devices would in principle allow any large group of strangers to coordinate on their very first attempt.
Once we dig deeper and identify which particular reaction function to play, it is no surprise that the best-reply reaction function stands out (i.e., to match the weakest link). In a sense, Theorems 5.1 and 5.2 together with the surrounding discussion make for a robustness exercise that verifies that, indeed, the intuitive best-reply reaction is the ideal reaction function to play in this case.
@@ 5.2 Public-good games
We turn to a different class of games for which, in contrast, best replies are ineffective. We have already seen that reaction functions introduce new ways for players to coordinate around Pareto-improving outcomes in strictly dominated actions (e.g. in the Prisoner’s dilemma). Now, we will further expand on this in the particular context of public-good games, in which each player’s investment benefits everybody, but the investment is a net loss to the player herself (e.g. Bergstrom et al., 1986). In public-good games, payoffs are given by
@@
with parameter $1/n<\lambda<1$. [^13] Outside this parameter range, high investment (for $\lambda\geq 1$) or zero investment (for $\lambda\leq 1/n$) is both safe, dominant, and efficient. High investment $(H,\dots,H)$ maximizes total payoffs $\sum_{j}u_{j}(a)$ but, in the strategic game, free-riding through $a_{i}=0$ is dominant. Hence, the best reply is simply to free-ride and never invest. Again, conventional safe play yields $\underaccent{\bar}{v}_{i}=u_{i}(0,\dots,0)=0$. Now, action profiles $a_{-i}$ and $a^{\prime}_{-i}$ are payoff equivalent if they add up to the same: $\sum_{j\neq i}a_{j}=\sum_{j\neq i}a^{\prime}_{j}$.
Whereas in the weakest-link games the key idea was to react to the minimum action of the others, in the public-good setting it is instead the average contribution that turns out to be essential. In particular, norm-proof and payoff-consistent reaction functions must match the average, either rounded up or down. [^14] For $z\in\mathbb{R}$, let $\lfloor z\rfloor,\lceil z\rceil\in\mathbb{Z}$ be the largest (smallest) integer no larger (smaller) than $z$.
@@
For public-good games, reaction function $R_{i}\in\mathcal{M}_{i}$ is norm-proof and payoff consistent if and only if there is a monotone function $f_{i}\colon\mathbb{R}\to\mathbb{Z}$ such that, for each $a_{-i}\in A_{-i}$,
$R_{i}(a_{-i})=f_{i}\left(\frac{1}{n-1}\textstyle\sum_{j\neq i}a_{j}\right)\in\left\{{\left\lfloor\frac{1}{n-1}\textstyle\sum_{j\neq i}a_{j}\right\rfloor},{\left\lceil\frac{1}{n-1}\textstyle\sum_{j\neq i}a_{j}\right\rceil}\right\}.$
Theorem 5.3 leaves two particularly focal candidates within the class of monotone, norm-proof, and payoff-consistent reactions: to match the average, either consistently rounding up or consistently rounding down. Of these, we will argue that rounding down (to be denoted $R^{*}_{i}$ below) fares better. Whereas rounding up is a more “generous” commitment, it does not effectively discipline free-riding, much like overly cooperative strategies do poorly in a repeated Prisoner’s dilemma (compare Axelrod, 1984). If all use reactions that round up, players would beneficially deviate by undercutting (for instance, by playing the constant reaction associated to investment H1H-1). For contrast, rounding down supports high investment in reaction-function equilibrium. On top of this, rounding down is safe across all parameter values, whereas “rounding up” is not for certain parameter values.
Beyond our equilibrium and refinement analysis, it is useful to benchmark reactions against a simple normative criterion: choose the investment that maximizes total welfare, subject to individual safe play. This also connects to the the experimental literature on social preferences (e.g. Charness and Rabin, 2002), where players sometimes behave as if they partially internalize others’ payoffs. Specifically, we define a reaction function $R_{i}\in\mathcal{R}_{i}$ to be welfare-maximizing conditional on safe play if, for each $a_{-i}\in A_{-i}$,
@@
As a parallel to Theorem 5.2, we define high-risk public-good games with parameters $\lambda$ such that $1/n<\lambda<H/(nH-1)$. The intuition is the same: there are still gains from coordinating on higher investment, but the returns are small compared to the risks involved in overinvesting relative to the others. For these games, matching the average, rounded down, is welfare-maximizing subject to safe play. [^15] For weakest-link games, this property instead singles out $\textit{BR}_{i}$. There is no reason for $i$ to exceed the minimum investment as this creates no added value to the others but only adds to $i$’s investment costs; it is also not beneficial for $i$ to undercut the minimum, because this harms all players more than it saves on $i$’s costs.
@@
For high-risk public-good games, $R_{i}\in{\cal M}_{i}$ is welfare-maximizing conditional on safe play if and only if
$R_{i}^{*}(a_{-i})=\left\lfloor\frac{1}{n-1}\textstyle\sum_{j\neq i}a_{j}\right\rfloor.$
In this way, reaction functions provide new ways to overcome tendencies to free-ride in public-goods games, which do not rely on, for instance, players forming sanctioning institutions (Kosfeld et al., 2009), being altruistic (Anderson et al., 1998), or the introduction of obligations (Galbiati and Vertova, 2008). Our main candidate $R^{*}_{i}$ is to match the average, rounded down, which appears to be a compelling rule-of-thumb for reaction-function play in public-good games. We do not claim that this reaction function always is the optimal strategic choice (e.g., it is better to free-ride if the others unconditionally contribute), but it appears always quite reasonable: it is simple, safe, and ensures that the player is no worse off than the average. If faced with a pool of free-riders, it specifies to free-ride as well; against both conditional and unconditional collaborators, high investment is a fixed point. In this way, it does reasonably well against most reaction functions. This is further corroborated by the simulation study in the Appendix. There, in a specific public-good setting, we find compelling evidence that a process of evolutionary selection would converge towards functions alike $R^{*}_{i}$.